nlewis7172:

The value of a professional basketball player's autograph rose 20% in the last year. It is now worth $288.00. What was it worth a year ago?

nlewis7172:

pls answer

jhonyy9:

let the value a year ago be x so given that after rose 20% the value is $288,00 x = ?

jhonyy9:

using the rule of 3 simple we can calcule the 20% what let be y x mean 100% so 1 y mean 20% so 0,2 y = 0,2*x so this mean that x +0,2x = 288 1,2x = 288 x = 288/1,2 = 240 hope this is understandably for you